OpenStudy (anonymous):

The Fibonacci sequence consists of the pattern 1,1,2,3,5,8,13,.... Look at the successive ratios of one term to the next. Make a conjecture.

OpenStudy (imstuck):

I will make the conjecture that you add the previous 2 terms together to get to the next one in the sequence. Check it...1+1 =2, the next term. 1+2=3, the next term. 2+3=5, the next term. And so on...

OpenStudy (anonymous):

Google what a conjecture is... I legitimately think that people no longer care to do any work themselves.

OpenStudy (anonymous):

Why are you on this website then?

OpenStudy (imstuck):

@adam.bloedel you seem angry. You don't have to help people here if you don't want to...no one is making you. Maybe some of these people are not as bright as you and need more explanation and patience.

OpenStudy (anonymous):

I'm not angry, I've literally watched people all day tell me they attempted it, when I know that they just don't want to do the work. Example, someone asked a question that I answered with a step by step for a much more complicated issue. Ten minutes later that same person asked another one, which I again step by step went through. Following that, the same question except I believe it was addition rather than subtraction was asked. That's called zero work ethic.

OpenStudy (imstuck):

I didn't read all your post...I didn't have to. Again, I say to you, if you don't want to be here listening to stupid people ask you to do their work for them, then go away. Not to be mean, I'm just saying that you could save yourself the frustration. My suggestion, apply at tutor.com and tutor people there...for money. It's truly satisfying, and the people who go there for help are paying very heavily to learn from us, like $40 an hour. So they are very wiilling to work and learn "how to do things for themselves".